# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

### v4

@@ Abstract
We estimate the number of street vendors in New York City. First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City. Second, we describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Third, we review ratio estimation and provide a theoretical justification based on the theory of point processes. Fourth, we use ratio estimation to calculate the total number of vendors, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,400 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. Finally, we evaluate the accuracy of the ratio estimator when the distribution of vendors is explained by a Poisson or Yule process, and we discuss several policy implications. In particular, our estimates suggest the American Community Survey misses the majority of New York City street vendors.
@@ 1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. [^1] Fees from mobile food vending licenses and permits provide nearly a million dollars in revenue to New York City each year (Mosher and Turnquist 2024). Fines paid by vendors provide approximately \$200,000 each year according to data from the New York City Office of Administrative Trials and Hearings. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
@@
In this paper, we propose an estimate for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in four sections. In Section 2, we review the process by which vendors receive licenses and permits, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We also derive standard errors.
In Section 4, we examine the validity of our assumptions. We evaluate whether the spatial distribution of respondents is representative by comparing it to the American Community Survey and administrative data, and we evaluate the Poisson process assumption by considering a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases. The Yule process may be appropriate if street vendors cluster within markets, in which case the data might have extra-Poisson variation. We rederive the standard errors from Section 3 to account for this variation.
In Section 5, we discuss some implications of our estimates. In particular, we find that though the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population. Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
@@ 2. Background and Survey Design
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and/or operate that unit (i.e., the vendors). Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm. [^2] For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees. Then Sally’s Salads counts as one business, two establishments, and seven vendors.
@@ 2.1 Background
A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available. Mobile food vending licenses are not limited. Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit. [^3] Note that our estimate reflects the number of street vendors as of 2021, prior to the implementation of Local Law 18 of 2021, which was enacted concurrently. Local Law 18 made several changes to the license and permit process. These changes do not impact our estimates and therefore are not discussed in this paper.
There are several types of mobile food vendor permits with varying restrictions. Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
@@ 2.2 Survey Design
The population of street vendors estimated in this paper is the population eligible for aid (as determined by SVP) that self-identifies as either a food or merchandise vendor. SVP found aid-eligible individuals through membership lists, referrals, and canvassing operations in which SVP affiliates visited neighborhoods. Survey operations continued until \$2,415,000 in funds were distributed, yielding 2,060 responses.
@@
The survey included 100 questions and was conducted in nine different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof. [^4] Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English. Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl and Quechua, typically spoke Spanish with a high level of fluency. The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status). Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
@@
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Section 7.1, Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Section 7.2, Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1. (For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
3. Methodology and Data Analysis
We estimate approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. We arrive at the estimate by using ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law.
In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership. In Section 3.2, we provide a theoretical justification of our estimates based on the theory of point processes, which, while more technical, is intended to provide a somewhat general argument for ratio estimation with spatial data when a design-based approach cannot be justified. Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 7.3. In Section 3.3, we apply the methodology to the survey data discussed in Section 2.2.
@@ 3.1 Ratio Estimation
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. We provide a simple explanation based on cross-multiplication, also called the rule of three. The purpose is to highlight the main assumption. A more rigorous derivation is provided in Section 3.2.
Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of type $i$. To fix ideas, let $A$ be New York City, let $i=0$ indicate mobile food vendors with a permit, and let $i=1$ indicate mobile food vendors without a permit. Also, let $n_{i}(A)$ denote the number of respondents of type $i$.
We observe $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$ from which we estimate $\theta(A)=\mathbb{E}\left[N_{0}(A)\right]$, the expected number of vendors without a permit. The expected number of mobile food vendors with or without a permit is then $\tau(A)=\theta(A)+N_{1}(A)$.
The ratio estimator of $\theta(A)$ is
$\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. (The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is accurate when the response rate, $p$, is approximately the same for each type, i.e.,
@@
in which case $\hat{\theta}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication.
The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit. If we further assume that the response rate does not vary by subregion, we can also estimate the total number of mobile food vendors by subregion. For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, e.g.,
@@
Then we can estimate $\theta(B)$ by the subregion estimator
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
where $\tilde{\theta}(B)$ is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
The subregion estimator is necessary when $N_{1}(B)$ is not observed and $\hat{\theta}(B)$ cannot be calculated directly. This is the case with the survey data described in Section 2.2. We observe $N_{1}(A)$, the number of permits city-wide (i.e., within region $A$), but not $N_{1}(B)$, the number of permits within subregion $B$. (We also observe $n_{0}(B)$ and $n_{1}(B)$.)
The subregion estimator $\tilde{\theta}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\theta}(B)$ with the estimate
$\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
The estimate $\tilde{N}_{1}(B)$ is also used to estimate the total number of mobile food vendors expected in subregion $B$. That is, we estimate $\tau(B)=\theta(B)+N_{1}(B)$ with
$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
@@ 3.2 Model
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to more closely examine the underlying assumptions and derive the standard errors of the estimates.
Let $\Pi_{i}$ denote an inhomogeneous spatial Poisson process referencing the location $x\in A\subset\mathbb{R}^{2}$ of each street vendor with status $i$ at the time the survey was conducted. As in Section 3.1, we fix ideas by letting $i=0$ indicate mobile food vendors with a permit and $i=1$ indicate mobile food vendors without a permit.
The Poisson process assumption may be justified by the law of rare events. New York City can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in any area is well approximated by a Poisson distribution. This approximation is accurate even if these probabilities are weakly dependent, see Freedman (1974). Other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
@@
In Section 3.1, we described the main assumption of ratio estimation: that the status and spatial distribution of the respondents is representative. We now make this statement precise. We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding mean measure. i.e.,
@@
so that by conditioning on $N_{1}(B)$, we arrive at the following probability model for the number of respondents with and without permits
$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\theta(B)\big{)}$
$n_{1}(B)\ |\ N_{1}(B)\sim$ $\ \text{Binomial}\big{(}p,\,N_{1}(B)\big{)}$
The maximum likelihood estimates for $p$ and $\theta(A)$ are
$\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}\qquad\text{and}\qquad\hat{\theta}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
Using increasing domain asymptotics, the estimator is asymptotically normal with mean $\theta(A)$ and standard error
$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{\dfrac{1}{p\,\theta(A)}+\dfrac{1-p}{p\,N_{1}(A)}}$
Substituting $\hat{p}$ and $\hat{\theta}(A)$ for $p$ and $\theta(A)$ yields the following plug-in estimator for the standard error,
$\hat{\text{SE}}[\hat{\theta}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
See Section 7.3.2 for details.
The maximum likelihood estimate for $\theta(B)$ is
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
which is asymptotically normal with mean $\theta(B)$ and standard error
$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}}$
The plug-in estimator for the standard error is
$\tilde{\text{SE}}[\tilde{\theta}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
To determine the subregion total $\tau(B)=\theta(B)+N_{1}(B)$, we estimate $\theta(B)$ and $N_{1}(B)$ with $\tilde{\theta}(B)$ and $\tilde{N}_{1}(B)=N_{1}(A)\,n_{1}(B)\,/\,n_{1}(A)$, yielding
$\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(B)+n_{1}(A\backslash B)}$
where $A\backslash B$ denotes the compliment of subregion $B$.

### v5

@@ Abstract
We estimate the number of street vendors in New York City. We first summarize the process by which vendors receive licenses and permits to legally operate in New York City. We then describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Finally, we calculate the total number of vendors using ratio estimation. We find that approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We also provide a theoretical justification of our estimates based on the theory of point processes and a discussion of their accuracy and implications. In particular, our estimates suggest the American Community Survey fails to cover the majority of New York City street vendors.
@@ 1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
@@
In this paper, we propose and justify an estimator for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in six sections:
In Section 2, we review the process by which vendors receive licenses and permits, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We then develop a theoretical justification based on the theory of point processes, which assumes the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes. Finally, we derive the standard errors.
In Section 4, we examine the sensitivity of the ratio estimator to deviations from the underlying assumptions. We first consider the case in which the spatial distribution of the survey respondents is not representative, and we examine a more general estimator in which sampling weights adjust for differential nonresponse. We then consider the case in which the Poisson process assumption is violated, and we examine a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases. The Yule process may be more appropriate if street vendors cluster within markets. In both cases, we rederive the standard errors.
In Section 5, we estimate the number of street vendors in New York City. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. We also find that this estimate is somewhat insensitive to deviations from our underlying assumptions. For example, we find that if the spatial distribution of survey respondents is not representative, the estimated number of street vendors in New York City could be as low as 21,000 and as high as 32,000.
In Section 6, we discuss some implications of our estimates. In particular, we find that although the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population. Our estimates suggest the American Community Survey misses the majority of New York City street vendors. This finding is important because the American Community Survey is commonly used for policy and advocacy, and our estimates shed light on the broader challenges of measuring the “gig economy.”
@@ 2. Background and Survey Design
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and/or operate that unit (i.e., the vendors). Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm. For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees. Then Sally’s Salads counts as one business, two establishments, and seven vendors.
@@ 2.1 Background
A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available. Mobile food vending licenses are not limited. Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.
There are several versions of a mobile food vendor permit with varying restrictions. Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
@@ 2.2 Survey Design
The population of street vendors estimated in this paper is the population eligible for aid (as determined by SVP) that self-identifies as either a food or merchandise vendor. SVP found aid-eligible individuals through membership lists, referrals, and canvassing operations in which SVP affiliates visited neighborhoods. Survey operations continued until \$2,415,000 in funds were exhausted, yielding 2,060 responses.
@@
The survey included 100 questions and was conducted in nine different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof. The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status). Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
@@
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1. (For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
3. Methodology
We estimate the number of street vendors that operate in New York City. We use ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law. In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership and highlights our main assumption. In Section 3.2, we develop a theoretical justification based on the theory of point processes, which, while more technical, allows us to explicitly state the estimands and underlying assumptions and to derive the standard errors. Inference is conducted using increasing domain asymptotics, and the details are outlined in Sections 7.1, 7.2, and 7.3.
@@ 3.1 Ratio Estimation
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. We provide a simple explanation based on cross-multiplication, also called the rule of three. The purpose is to highlight the main assumption. A more rigorous derivation is presented in Section 3.2.
Consider a fixed region $A$, and let $N_{i}(A)$ denote the number of individuals in region $A$ of status $i$. To fix ideas, let $A$ be New York City, let $i=0$ indicate mobile food vendors without a permit, and let $i=1$ indicate mobile food vendors with a permit. Also, let $n_{i}(A)$ denote the number of respondents of status $i$.
We observe $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$ from which we would like to determine $N_{0}(A)$, the number of vendors without a permit. The total number of mobile food vendors with or without a permit is then $N_{0}(A)+N_{1}(A)$.
We define the ratio estimator as
$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. (The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is approximately equal to $N_{0}(A)$ when the response rate, $p$, is approximately the same for each status, i.e.,
@@
in which case $\hat{\Lambda}_{0}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication. It follows that $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$ is approximately equal to the total $N_{0}(A)+N_{1}(A)$.
The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit. If we further assume that the response rate does not vary by subregion, we can also determine the total number of mobile food vendors by subregion. For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, i.e.,
@@
Then the subregion estimator
$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
is the solution for $N_{0}(B)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(B)$.
The subregion estimator is useful when $N_{1}(B)$ is not observed and $\hat{\Lambda}_{0}(B)$ cannot be calculated directly. This is the case with the survey data described in Section 2.2. We observe $N_{1}(A)$, the number of permits city-wide, but not $N_{1}(B)$, the number of permits within subregion $B$. (Recall that the vending locations of the respondents are recorded and thus $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subset A$.)
The subregion estimator $\tilde{\Lambda}_{0}(B)$ works by replacing the term $N_{1}(B)$ in the ratio estimator $\hat{\Lambda}_{0}(B)$ with
$\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}$
which is justified when
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{1}(B)}{N_{1}(B)}$
This suggests the estimator
$\tilde{\tau}(B)=\tilde{\Lambda}_{0}(B)+\dfrac{N_{1}(A)\,n_{1}(B)}{n_{1}(A)}=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
is approximately equal to the subregion total $N_{0}(B)+N_{1}(B)$.
@@ 3.2 Model
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to explicitly state the estimands and underlying assumptions and to derive the standard errors.
Let $\Pi_{i}$ denote an inhomogeneous Poisson process referencing the location $x\in A\subset\mathbb{R}^{2}$ of each vendor with status $i$ at the time the survey was conducted. As in Section 3.1, we fix ideas by letting $i=0$ indicate mobile food vendors without a permit and $i=1$ indicate mobile food vendors with a permit.
The Poisson process assumption may be justified by the law of rare events. Any sufficiently large area can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in that area is well approximated by a Poisson distribution. This approximation is accurate even under weak dependence (Freedman 1974). Other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
@@
The data are the locations of the respondents by status (i.e., the realizations of $\pi_{i}$) and the number of vendors in region $A$ with a permit, $N_{1}(A)$. In other words, we assume $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$, while $N_{0}(B)$ and $N_{1}(B)$ are unobserved with the exception of $N_{1}(A)$. Our goal is to estimate the expected number of street vendors in subregion $B\subseteq A$ without a permit,
@@
and the expected total with or without a permit,
$\tau(B)=\Lambda_{0}(B)+\mathbb{E}\left[N_{1}(B)\,|\,N_{1}(A)\right]$
We condition on $N_{1}(A)$ to reflect the fact that while the number of vendors of either status in any subregion $B\subsetneq A$ is random, the total number of vendors with permits citywide is fixed by law.
In Section 3.1, we described the main assumption of ratio estimation: that the status and spatial distribution of the respondents is representative. We now make this statement precise. We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding intensity. i.e.,
@@
so that by conditioning on $N_{1}(A)$, we arrive at the following probability model for the number of respondents with and without permits
$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\Lambda_{0}(B)\big{)}$
$n_{1}(B)\,|\,N_{1}(A)\sim$ $\ \text{Binomial}\big{(}N_{1}(B),\,p\,q(B)\big{)}$
where
$q(B)=\frac{\Lambda_{1}(B)}{\Lambda_{1}(A)}$
The maximum likelihood estimates for $\Lambda_{0}(A)$ and $p$ are
$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\qquad\text{and}\qquad\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}$
Using increasing domain asymptotics, $\hat{\Lambda}_{0}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
The maximum likelihood estimate for the total $\tau(A)$ is $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$, which has the same standard error as $\hat{\Lambda}_{0}(A)$. See Section 7.1.2 for details.
The maximum likelihood estimate for $\Lambda_{0}(B)$ is
$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
which is asymptotically normal with mean $\Lambda_{0}(B)$ and standard error
$\tilde{\text{SE}}[\tilde{\Lambda}_{0}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
Finally, the maximum likelihood estimate for the subregion total $\tau(B)$ is
$\tilde{\tau}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$